OpenStudy (anonymous):

a student received scores of 88,82, and 84 on 3 quizzes. if tests count twice as much as quizzes, what is the lowest score the student can get on a test to achieve an average of at least 70?

OpenStudy (anonymous):

take weighted average and set it equal to 70. that will give you the minimum test score required,

OpenStudy (amistre64):

hmmm.... yeah, weight them. the value of a quiz is worth 1/3 and the value of a test is worth 2/3 88+82+84 ---------- * 1/3 + testgrade(2/3) = 70, if i see it right 3

OpenStudy (amistre64):

or in one fell swoop: testgrade = (70 - (88+82+84)/3) / (2/3) = -22 ? gonna have to rethink that maybe

OpenStudy (amistre64):

i assumed 1 test since it said "a test" :)

OpenStudy (amistre64):

ack, i think i left out the "1/3" in my fell swooper lol

OpenStudy (amistre64):

testgrade = (70 - (254/3)(1/3)) / (2/3) = 62' 2/3